Bally Manufacturing sent Intel Corporation an invoice for machinery with a $14,200 list price. Bally dated the invoice August 04 with 5/10 EOM terms. Intel receives a 40% trade discount. Intel pays the invoice on August 17. What does Intel pay Bally?
Answer #1

Net payable = 14,200 - Trade discount 40%

= 14,200 - 40%

= 8520

Intel pays to Bally = 8520 - (5% * 8520)

= 8520 - 426

= 8094

Comment if you face any issues